### v2

# INQUIRE-Search: A Framework for Interactive Discovery in Large-Scale Biodiversity Databases

Abstract
Large community science platforms such as iNaturalist contain hundreds of millions of biodiversity images that often capture ecological context on behaviors, interactions, phenology, and habitat. Yet most ecological workflows rely on metadata filtering or manual inspection, leaving this secondary information inaccessible at scale. We introduce INQUIRE-Search, an open-source system [^1] The codebase will be published open-source at https://github.com/Beery-Lab/INQUIRE-Search/tree/main. that enables scientists to rapidly and interactively search within an ecological image database for specific concepts using natural language, verify and export relevant observations, and utilize this discovered data for novel scientific analysis. Compared to traditional methods, INQUIRE-Search takes a fraction of the time, opening up new possibilities for scientific questions that can be explored. Through five case studies, we show the diversity of scientific applications that a tool like INQUIRE-Search can support, from seasonal variation in behavior across species to forest regrowth after wildfires. These examples demonstrate a new paradigm for interactive, efficient, and scalable scientific discovery that can begin to unlock previously inaccessible scientific value in large-scale biodiversity datasets. Finally, we emphasize using such AI-enabled discovery tools for science call for experts to reframe the priorities of the scientific process and develop novel methods for experiment design, data collection, survey effort, and uncertainty analysis.
1 Introduction
1.1 The Need for Tools to Make Ecological Databases Discoverable
Image data have become a key component of ecological research, collected to document and analyze biodiversity, species interactions, and environmental changes across spatial and temporal scales [oliver2023camera, kays2020born]. The widespread adoption of digital photography and the deployment of automated sensors has led to a significant increase in the size and scope of ecological image databases. As an example, community science platforms like iNaturalist [iNaturalist2024] contain hundreds of millions of image-based species observations, an unprecedented scale, with tremendous potential for ecological science [callaghan2021three].
So far, the scientific use of such databases has primarily focused on documenting and analyzing species occurrences (the identification of a species at a specific place and time) [kumar2019using, humphreys2019seasonal]. But many facets of ecological study, such as biotic and abiotic interactions, resource use, and habitat associations, require information that extends beyond species identification and location alone. We refer to this contextual information as secondary data: ecologically relevant features that extend beyond the primary observation [pernat2024overcoming, marques2024retrieving, davison2025automated]. In image data, each observation shows organisms in their realized microhabitats, often with co-occurring species, observable behaviors, or signs of disturbance. Efficiently unlocking this information would enable these large image databases to realize their potential scientific utility. [beccacece2025exploring, pernat2024extracting, hu2025global].
[…]
Figure 1: INQUIRE-Search pipeline overview. Once a scientific question is posed, a user can leverage our pipeline to flexibly discover the relevant data and perform analysis iteratively. Our pipeline involves the following stages: (1) Query and Prioritize to retrieve relevant images from the iNaturalist repository using natural language queries, (2) Refine and verify to manually allocate human effort and inspect top retrievals for correctness, and (3) Export and analyze to perform downstream analysis on the discovered data. This system is intended to be iterative: users can return to (1) after (2) or even (3) and repeat the process until sufficient data collection has been fulfilled.
1.2 Utilizing Machine Learning to Process Large Scale Data
Recent advances in machine learning, particularly in computer vision, provides valuable tools for dataset curation and filtering. In ecology, automated species detection and classification models are used at scale to filter large-scale databases [norouzzadeh2018automatically, willi2019identifying]. However, the supervised learning approaches to train these models require large, well-labeled datasets, which are often scarce and expensive to annotate for ecological information of interest. Moreover, scientific discovery often targets rare or context-dependent phenomena—such as predator-prey interactions, behavioral displays, or signs of disturbance—that cannot be easily captured through a fixed set of pre-defined categories. For example, modeling interactions may require defining $\binom{n}{2}$ species-pair categories for each interaction type, quickly leading to an impractically large and sparsely populated label space. This traditional closed-set design limits conventional supervised models, constraining their ability to generalize beyond predefined categories and hindering their use in open-ended ecological exploration.
Vision-language models (VLMs) offer a powerful alternative. VLMs are trained on large-scale image-text datasets, creating a joint embedding space that supports similarity-based search across modalities. Crucially, VLMs enable natural language-based querying of image databases, allowing researchers to search and analyze ecological imagery through flexible, human-readable prompts. Some of these models exhibit broad, open-world understanding of visual content and can interpret complex ecological scenes without task-specific supervision [vendrow2024inquire].
The cross-modal capabilities of VLMs emerge from their training objective, which encourages the model to map both images and language into a single high-dimensional "representation space" where similar concepts are close together, and different concepts are far apart [radford2021learning]. As a simple example, an image of a hermit crab wearing plastic as a shell would be embedded into this space closer to the corresponding textual description of the scene (i.e. "a hermit crab with plastic trash as a shell"), and farther from the embedding of the text "a yellow car," as seen in Figure˜2. These embeddings capture the semantic content of images and text across many different dimensions—color, texture, shape, taxonomy, phylogeny, habitat, etc. These models therefore enable us to search for relevant images in a database through a simple similarity search—ranking each image in the database based on its distance to the embedding of a text query.
Figure 2: (Top) Joint vision-language representation space. VLMs are trained to learn a joint embedding space between vision and text modalities, where similar semantic concepts are stored closer together and different concepts are farther apart. (Bottom) Image-text search using FAISS indices. At search-time, the input text query is embedded by the text encoder of the VLM. Then the query embedding vector is compared against the pre-computed iNaturalist image embeddings to retrieve the most relevant images.
This type of embedding-enabled search with VLMs has been explored for ecological image databases in several recent works. WildCLIP [gabeff2024wildclip] fine-tuned a VLM with camera trap data and ecology-specific language descriptions. In training over large quantities of data, this model was able to generalize from concepts seen during training to search large wildlife datasets for new behaviors or habitat details. More recently, the INQUIRE Benchmark (INQUIRE-Bench) [vendrow2024inquire] was introduced to evaluate state-of-the-art VLMs with expert-level open-ended ecological queries across five million iNaturalist 2024 images. While the benchmark continues to pose a significant challenge for even the most capable vision-language models, it was instrumental in identifying which vision-language models perform best for these types of ecological questions. Both works focus primarily on benchmarking and improving retrieval accuracy, and clearly demonstrate the potential of such an approach to unlocking new knowledge in databases. Realizing this potential, however, requires a system that facilitates rigorous ecological inference by enabling scientists to rethink experimental design and carefully define data discovery and analysis workflows for addressing specific scientific questions.
1.3 INQUIRE-Search: a System for Interactive and Efficient Ecological Retrieval
To address this need, we introduce INQUIRE-Search, a system that enables efficient, expert-driven, iterative retrieval from large ecological image datasets. The system combines the open-world understanding of modern VLMs with question-driven methodology of ecological research: INQUIRE-Search treats large-scale dataset exploration as a process of asking, refining, and validating questions. This AI-assisted approach shifts the methodological focus from exhaustive manual labeling to strategic query design and iterative refinement, guided by both human expertise and AI retrieval results.
INQUIRE-Search is built on the principles of natural language interaction, expert-guided refinement, and efficiency. Together, these principles transform static image repositories into interactive engines for discovery. The system implements this approach through an interactive, three-stage workflow illustrated in Figure˜1:
  • Query and Prioritize: Researchers pose ecological questions in plain language (e.g. "California condor with a green "26" tag on its wing"). INQUIRE-Search retrieves and prioritizes data, ranking images by their semantic relevance to the query.
  • Refine and Verify: Scientists rapidly assess and confirm ecological validity of candidate results. This verification is flexible, allowing researchers to decide how much manual labor to allocate. Based on the results, they can also iterate by manually refining or reformulating their query to sharpen the search and ensure sufficient data collection.
  • Export and Analyze: Curated image sets can be exported with associated metadata for quantitative or qualitative analysis, supporting downstream studies or dataset curation.
INQUIRE-Search enables users to search hundreds of millions of natural-world images using natural language queries, efficiently retrieving relevant observations to minimize manual browsing. To summarize, INQUIRE-Search combines (1) a state-of-the-art vision-language embedding model that maps images and text into a shared semantic space, (2) a high-performance vector index that enables fast similarity search across hundreds of millions of embeddings, and (3) a lightweight, browser-based interface that supports filtering, inspection, verification, and data export to achieve this. Together, these components allow ecologists to iteratively query, refine, and curate datasets in real-time, transforming unstructured community science imagery into analyzable evidence for ecological research. We provide a technical overview of the pipeline and specific implementation details in the Appendix.
2 Case Studies
We worked with a diverse set of ecological experts to define a series of scientific questions that would potentially be addressable with information hidden within the iNaturalist database. We worked together to use INQUIRE-Search to address each of these questions, designing each experiment to test both the usability and utility of the tool and test the limits of what insights the system could help uncover. Each case study explores a distinct ecological question: (1) seasonal variation in bird diets, (2) post-fire forest regeneration, (3) spatio-temporal patterns of wildlife mortality, (4) plant phenology across seasonal cycles, and (5) individual Re-ID in humpback whales. These examples show how natural language-guided image search and analysis can provide a scalable and flexible approach for extracting ecological insights from large, unstructured image datasets.
2.1 Seasonal Variation in Bird Diets
2.1.1 Task Description
Diet plays a pivotal role in vertebrate life history, fundamentally influencing species’ survival, recruitment, and overall fitness. As a response trait, it is shaped by long-term environmental changes, with shifts in climate and resource availability driving dietary diversification across a wide range of taxa [burin2016omnivory, sibly2012energetics, belmaker2012global]. As an effect trait, diet provides insight into an organism’s ecological impact and its biotic interactions, reflected in its resource use and trophic position. Because of this dual role, diet has been central to advancing our understanding of key ecological and evolutionary processes, including species coexistence [abrahamczyk2015morphological], character displacement and evolutionary divergence [grant2006evolution], vertebrate diversification [burin2016omnivory, price2012tempo, rojas2012role, felice2019dietary], community assembly [lu2023food], and the impacts of global change on biodiversity [ko2014dietary].
Figure 3: Avian diets vary by species and time of year. Understanding diet enables us to connect seasonal resource use with the ecological processes that shape species interactions and community structure.
Although dietary databases exist for some vertebrate groups [wilman2014eltontraits]—including those capturing seasonal variation in diet [murphy2023SAviTraits]—they are typically compiled from published literature and often lack complete information for many species. For instance, SAviTraits 1.0 [murphy2023SAviTraits] provides seasonal dietary data for over 10,000 extant bird species, yet only about 10% of those species are recorded as exhibiting seasonal dietary variation. This might reflect not true biological invariance, but rather the absence of data on temporal shifts in diet for most species. Combined with known uncertainties in dietary classifications, it is clear that many species remain poorly studied—some may indeed exhibit seasonal dietary changes that simply have not been documented in the literature. Addressing these data gaps through traditional field-based methods would be prohibitively labor-intensive and infeasible at broad spatial scales. Current search and metadata-based filtering tools on iNaturalist are insufficient for comprehensive data extraction. For example, among more than 23,000 relevant observations, only 44 are annotated with the tag "feeding" and none are labeled "feeding on fruit." Consequently, identifying relevant cases would require manual inspection of all 23,000+ observations, which is prohibitively inefficient.
In this case study, we investigate whether INQUIRE-Search could enable systematic analysis of seasonal dietary variation in birds by searching through millions of community-sourced photographs on iNaturalist. To validate INQUIRE-Search’s performance, we compare our results against dietary information captured in SAviTraits [murphy2023SAviTraits]. We focus on five bird species with different but known diets: Gray-cheeked Thrush (Catharus minimus), Ancient Murrelet (Synthliboramphus antiquus), American Tree Sparrow (Spizelloides arborea), Red-bellied Woodpecker (Melanerpes carolinus), and American Robin (Turdus migratorius) as shown in 3. In SAviTraits, the certainty designation—derived from the number of citations supporting a species’ dietary description and the length of that description—exceeds 0.85 for these species, indicating they are among the best-studied birds in terms of diet. This makes both SAviTraits and these particular species a comprehensive and robust reference for comparison with INQUIRE-Search.
2.1.2 Methodology
We devised a systematic approach using INQUIRE-Search to extract seasonal dietary information:
  1. 1.
    For each target species, we construct natural language search queries specifying the bird species and diet type (e.g., "American robin with invertebrate in its mouth"), using the dietary classification system used in SAviTraits. This classification system includes 7 dietary categories: (1) invertebrate, (2) vertebrate, (3) seed, (4) fruit, (5) nectar and pollen, (6) carrion or animal decay, and (7) other plant matter. We applied taxonomic and temporal filters to restrict results to the target species and season (June-August for boreal summer and December-February for boreal winter), respectively. For each species, we ran 14 searches: one for each diet type-season combination.
  2. 2.
    Two researchers independently searched for and analyzed 500 images per species-diet type-season combination, each using the INQUIRE-Search tool. For each image, the researchers marked whether it contained clear dietary information for the food item corresponding to the query (i.e., the food item is visible and recognizable in the bird’s mouth).
  3. 3.
    We counted the number of marked images for each species, diet type, and season. INQUIRE-Search based dietary information was then derived from the proportion of marked images of each food type.
Figure 4: (Top) Query & Verification. Top INQUIRE-Search returns for "Red-bellied Woodpecker" and "Invertebrate" with incorrect or irrelevant images removed through expert review. (Bottom) Diet comparisons. INQUIRE-Search yields dietary patterns that align closely with those documented in the SAviTraits reference dataset.
2.1.3 Results and Analysis
Our analysis revealed substantial variation in the availability of dietary information across the five study species 4. American Robin yielded the highest number of informative (i.e., containing dietary information) images (669 total), followed by Red-bellied Woodpecker and American Tree Sparrow (395 and 188, respectively), while for both Ancient Murrelet and Gray-cheeked Thrush we found a negligible number of images containing dietary information. Among the species for which dietary information was available, seasonal patterns in image availability were also evident: American Robin had more informative images in summer than winter, while Red-bellied Woodpecker and American Tree Sparrow had more images available for winter months.
For the three species with sufficient data, dietary composition derived from INQUIRE-Search closely matched reference data from SAviTraits Figure˜4. For example, in summer, American Robin exhibited an even split between fruit and invertebrate consumption (50% each) in both datasets, while Red-bellied Woodpecker showed a diet dominated by seeds (50% in INQUIRE-Search vs. 40% in SAviTraits), with smaller contributions from invertebrates, fruits, and plant material. In winter, all three species exhibited diets strongly dominated by a single food type, and we again observed close agreement between INQUIRE-Search outputs and SAviTraits data.
2.1.4 Findings
We demonstrate that our method can effectively reproduce dietary patterns documented in established literature, provided that sufficient image data are available. The close alignment between INQUIRE-Search-derived avian diets and SAviTraits reference data highlights the potential of community science imagery to yield accurate ecological insights when properly analyzed through a systematic search methodology. However, this approach is most effective for commonly photographed species such as American Robin and Red-bellied Woodpecker. Two of the species we examined lacked sufficient image data for meaningful analysis—despite being well-studied in the scientific literature and represented in databases like SAviTraits—underscoring a fundamental constraint of community-sourced data. We anticipate this limitation will be even more pronounced for species that are genuinely understudied in the literature and for which new dietary information is most needed, as these species are also likely to be underrepresented in iNaturalist, and thus not retrievable with INQUIRE-Search. Overall, this case study shows INQUIRE-Search’s utility for rapid, large-scale analysis of visual biodiversity data, while also emphasizing its current constraints.
2.2 Post-fire Forest Regrowth
2.2.1 Task Description
Understanding how forests recover following fire is crucial for monitoring and predicting long-term ecosystem resilience. In the western US, wildfire is and has been the dominant disturbance for millennia [whitlock2008long, marlon2012long]. As a result, forests in the region are largely fire adapted and depend on wildfire for reproduction, competition reduction, and nutrient cycling [keeley2011fire]. However, the long-term consequences of fire suppression paired with recent warming and drying trends resulting from climate change are leading to wildfires that are more frequent, severe, and intense [abatzoglou2016impact, juang2022rapid]. These shifts are exceeding the adaptive capacity of many forest systems and jeopardizing their long-term resilience [harvey2016burn, davis2023reduced].
Despite the ecological importance of post-fire recovery, quantitatively assessing forest regrowth remains limited with existing survey approaches. Satellite-derived data provides broad coverage, but is generally available at spatial scales too coarse to detect individual trees and the spectral resolution of most satellites is not high enough to differentiate species [xu2021spatial]. Additionally, young trees are often spectrally indistinguishable from herbaceous species for up to 20 years after a fire [kiel2022trees]. Thus, existing satellite-based approaches struggle to capture early post-fire successional dynamics.
Field-based approaches provide valuable insights as tree seedlings are measured directly, but field campaigns are often limited in spatial coverage and temporal frequency due to cost and effort constraints. iNaturalist is a burgeoning data source that could help to overcome these constraints by providing geotagged images sourced by citizen scientists. INQUIRE-Search’s ability to efficiently search and classify iNaturalist images has the potential to reveal post-fire species and vegetation communities that are otherwise invisible to traditional remote-sensing and field-based approaches alike.
Figure 5: (Left) Forest regrowth following the 2012 High Park Fire in Colorado. With wildfires becoming more frequent and severe, fine-scale characterization of vegetation recovery is critical for understanding forest resilience. (Right) MTBS burn severity categories. We aggregate the burn categories from 30m Landsat resolution to a spatial resolution matching iNaturalist metadata (0.01 degree, $\sim$1.1km)
In this case study, we use INQUIRE-Search to identify and classify emergent forest communities growing after the 2012 High Park fire in Colorado (Figure˜5). We use INQUIRE-Search to search through nearly 20,000 ground-level iNaturalist images taken within the fire perimeter and provide a set of images containing individuals representing two target forest types: coniferous and deciduous. To test how burn severity impacts forest regeneration outcomes, we also examine the relationships between burn severity categories as classified by the Monitoring Trends in Burn Severity (MTBS) dataset (right side of Figure˜5) and each forest type.
2.2.2 Methodology
We designed a systematic approach using INQUIRE-Search to identify individuals representing emergent post-fire forest communities for each burn severity category:
  1. 1.
    To restrict images to a spatial subset containing the burn area, we bounded our search by latitude (40.57 to 40.75 degrees North) and longitude (105.18 to 105.54 degrees West) (Figure˜5). We also manually checked image timestamps to ensure all images were captured after the High Park fire.
  2. 2.
    For each target forest type, we constructed a natural language search specifying the forest type and restricting images to areas with evidence of past burning. The specific search terms are as follows:
    • “young coniferous trees in burned forest”
    • “young deciduous trees in burned forest”
  3. 3.
    We exported and analyzed 200 images for each search. Using reported image coordinates, we restricted images to burned areas within the High Park MTBS fire perimeter using the sf package in R3 [pebesma2018simple].
  4. 4.
    For each image, a researcher marked informative images containing the forest community of interest (i.e., an individual of the target forest type could be clearly distinguished). We allowed individual images to be marked as informative in multiple searches if the image contained individuals representing both forest types (e.g. a coniferous and deciduous tree seedling). While processing images for searches related to trees, we also decided to include trees and shrubs within the same target forest type as it was difficult to distinguish young trees from shrubs in images.
  5. 5.
    We extracted MTBS burn severity categories for the selected image locations and related burn severity to regeneration outcomes. Burn severity categories were aggregated from the native 30-m Landsat resolution to 0.01 degree resolution (Figure˜5) or consistency with the exported image coordinates. We aggregated burn severity categories using the most common burn severity category among contributing pixels using the terra package in R4 [hijmans2022package].
  6. 6.
    We count the number of marked images for each forest type and burn severity category.
2.2.3 Results and Analysis
Figure 6: (Top) Query & Verification. Top INQUIRE-Search returns for "Young coniferous trees in burned forest" with incorrect or irrelevant images removed through expert review. The images were marked correct if the right type of tree is visible in any part of the image, even if not in the foreground. (Bottom) Tree observations across burn severity. Distribution of observed young coniferous and deciduous trees in community science collected images show strong relationships with burn severity regions.
Of the 200 images exported from INQUIRE-Search for each search, there were 100 (coniferous) and 112 (deciduous) images with reported coordinates within the High Park burn perimeter. We marked more informative images for deciduous forest species (78) than coniferous (45). We also marked five images depicting both coniferous and deciduous individuals, indicating areas that could recover into mixed forest communities
Among images marked for analysis, we found strong relationships between burn severity and post-fire regeneration of both coniferous and deciduous forest types. For both forest types, most marked images of regenerating vegetation occurred in areas classified as unburned-to-low or low burn severity, with few observations in moderate or high severity areas (Figure˜6, left). However, coniferous regeneration showed a stronger negative association with increasing burn severity: 93.3% of coniferous images were located in unburned-to-low or low severity areas, with only 4.4% and 2.2% in moderate and high burn severity areas, respectively (Figure˜6, right). In contrast, deciduous regeneration was somewhat more evenly distributed, with 84.6% of marked images in unburned-to-low or low severity areas, and a higher proportion (7.6% each) in areas burned at moderate and high severity (Figure˜6, right). These patterns suggest that post-fire forest regeneration—or at least regeneration captured by citizen images—is more common in lower severity areas, with coniferous species showing greater sensitivity than deciduous species.
2.2.4 Findings
In this case study, we demonstrate that INQUIRE-Search can be used to effectively identify post-fire forest regeneration patterns across burn severity gradients. Our approach successfully distinguished between regenerating coniferous and deciduous forest types and provided spatially explicit insights into post-disturbance recovery within the first decade after the High Park fire. We found that regeneration was most commonly observed in unburned areas or areas burned at low severity, with coniferous species showing greater sensitivity to increasing burn severity than deciduous species. This is consistent with previous research documenting that increased fire severity decreases conifer regeneration likelihood both regionally and specifically after the High Park fire [davis2023reduced, wright2017post]. Overall, this method offers a scalable and accessible tool for monitoring ecological recovery using public image databases.
Despite INQUIRE-Search’s utility for harnessing citizen science data and identifying post-fire vegetation communities, several limitations remain. Image coordinates are rounded to 0.01° to match the limiting spatial precision of many iNaturalist records, but this prevents fine-scale matching to burn severity or site attributes such as seed source or soil moisture. Classification from user-submitted photographs also introduces uncertainty: landscape images often obscure species identification, and deciduous seedlings, saplings, shrubs, and herbaceous plants can appear morphologically similar. Our expert reviewer frequently could not distinguish these groups, reflecting a broader challenge shared with remote-sensing approaches: early-seral vegetation often remains visually and spectrally ambiguous for years after fire. Finally, iNaturalist data are opportunistic and biased toward accessible, aesthetically appealing areas, which may result in severely burned sites being underrepresented (Fig. 3). Thus, while INQUIRE is a powerful supplemental tool for assessing post-fire recovery, its results should be interpreted alongside field data and remote-sensing derived products to minimize bias.
2.3 Wildlife Mortality
Figure 7: (Left) Wildlife mortality trends vary by season and location. Collisions with human-made structures such as windows are a major cause of bird deaths, varying seasonally and with urbanization. (Right) Comparing avian mortality in rural vs urban regions. Using INQUIRE-Search, we compared image-based evidence of avian mortality between urban (Boston, MA) and rural (Pioneer Valley, MA) sites to examine how seasonal risks differ across anthropogenic contexts.
2.3.1 Task Description
Understanding patterns of avian mortality is fundamental to conservation efforts, particularly given the dramatic decline of nearly 3 billion birds in North America since 1970 [rosenberg2019decline]. Mortality risk varies seasonally as birds encounter different threats across the annual cycle [loss2015direct, marra2015call], and these patterns differ across landscapes due to variation in local hazards and bird abundance. Collisions with buildings—one of the leading human-driven causes of bird death [loss2014estimation]—are strongly seasonal, peaking during migration [riding2021multi, scott2023causes], and are further shaped by anthropogenic context, with tall structures in rural areas often generating disproportionately high collision rates [hager2017continent]. Understanding how mortality patterns shift across space and time is therefore essential for designing targeted interventions and addressing ongoing population declines.
However, documenting spatio-temporal patterns of animal mortality presents significant methodological challenges. Direct observation of dead animals in the wild is relatively rare because mortality events occur infrequently and alternative approaches such as marking or tracking individuals are labor-intensive, expensive, and only feasible for certain taxa at limited scales [yanco2025tracking]. Leveraging large-scale image datasets to identify where and when animals die would fill key gaps in our ability to study the patterns and processes of wildlife mortality.
In this case study, we used INQUIRE-Search to investigate seasonal avian mortality dynamics in an urban (Boston, MA) and rural (Pioneer Valley, near Amherst, MA) setting and compare relative mortality dynamics between the two sites. We hypothesized that (1) mortality rates would be higher in urban contexts, due to the greater density of man-made infrastructure that poses a collision risk [loss2015direct] and (2) mortality rates would be greatest during seasonal migrations when avian densities are highest [scott2023causes].
2.3.2 Methodology
We used a systematic approach to estimate mortality rates at both sites, by querying INQUIRE-Search:
  1. 1.
    We queried INQUIRE-Search for "dead bird" in each of the two equal-sized bounding boxes. We visually inspected results and removed records of live birds, single feathers, non-avian species, and duplicate observations of the same mortality event (duplicates defined by shared species, month, latitude, and longitude).
  2. 2.
    To compare the number of observed mortalities to the overall number of iNaturalist observations, we also downloaded the total number of iNaturalist observations of birds within those spatial bounding boxes using the rinat package in R [barve2022rinat].
  3. 3.
    We summarized both mortalities and observations by month for each study area and then calculated monthly mortality "rate" as the monthly mortality count divided by monthly observation count. Because we expect the baseline bird abundance, observation effort, and mortality rates to differ among sites, we normalized each month’s effort by dividing it by the mean monthly site-specific mortality rate and then taking the $\log_{2}$. This results in a "Mortality Index" which controls for differences in observer effort and underlying bird population abundances. Taking the log of mortality rate symmetrically centers the index on zero such that values of this index above 0 represent the fold-change above mean mortality rates whereas values below 0 reflect the converse.
Figure 8: Seasonal patterns in relative bird mortality for an urban (Boston, blue) and rural (Pioneer Valley, orange) region of Massachusetts. Mortality is shown as a $\log_{2}$ index scaled to each site’s annual mean (0), with units representing fold-change. Positive values indicate above-average mortality, negative values below.
2.3.3 Results and Analysis
The results show distinct patterns between urban and rural settings. Boston exhibits notable peaks in mortality during the spring and fall, moderate mortality rates during the summer, and low mortality rates in winter. In contrast, the Pioneer Valley shows elevated mortality rates in summer and fall, moderate mortality rates in early winter, and low mortality rates in late winter and spring. The high mortality rates at both sites in September and October are consistent with expectations of peak migratory passage—and the concomitant risk of building collisions—being highest in the fall in the eastern United States [horton2019bright]. Interestingly, Boston exhibited higher mortality rates during spring migration which may reflect differential migratory passage rates or other differences in site-specific risk factors [scott2023causes]. Finally, the Pioneer Valley did not exhibit the winter drop in mortality rates that we observed in Boston. While this could reflect differences in the avian communities or risks between the two sites, we cannot rule out the possible role of increased detection probability due to reduced vegetative cover disproportionately affecting the more rural Pioneer Valley.
2.3.4 Findings
We demonstrate that INQUIRE-Search can detect distinct spatiotemporal patterns in avian mortality through analysis of community-sourced photographs. The contrasting seasonal dynamics between urban (Boston) and rural (Pioneer Valley) sites provide a compelling demonstration of the importance of anthropogenic context in driving avian demographics.
However, several important limitations must be considered when interpreting these results. First, detection biases likely vary substantially between urban and rural environments. In urban settings, dead birds are more likely to be discovered on sidewalks, building perimeters, and other high-visibility locations where citizen scientists frequently travel. On the other hand, rural mortality events may be significantly underdetected due to the tendency for carcasses to occur in forested or remote areas where they are quickly removed by scavengers or decompose unnoticed. Second, the observer demographics and motivations for using iNaturalist may differ between urban and rural areas, potentially introducing additional sampling biases. Third, our mortality index assumes that the ratio of dead bird observations to total bird observations provides a meaningful proxy for actual mortality rates, but this assumption may not hold if detection probabilities vary systematically across seasons, locations, or species.
2.4 Resolving Plant Phenology
2.4.1 Task Description
Figure 9: Four pheonological stages of common milkweed. Example observations identified using INQUIRE-Search that correspond to the four phenological stages analyzed in this study: emergence, flowering, seeding, and senescence.
Studying how the timing of specific life history events is changing is a primary way researchers are detecting biodiversity response to ongoing global change [parmesan2003globally]. Often highly malleable, phenological shifts in the timing of events like plant germination and reproduction, can serve as early warning signs that precede other biodiversity responses like demographic, genetic, or geographic shifts which can threaten the persistence of communities.
Traditionally, assessing phenological changes involves a significant amount of field study, to record and track the timing of specific events year after year. For example, Austin et al. (2024) [austin2024climate] use multiple phenology datasets gathered over 5 years comprising weekly transects during growing months (March-November) to understand shifts in flowering onset and duration in a single plot in the Ozark Highlands in Missouri, USA. While these sorts of datasets permit fine scale assessments of how species in a specific community are responding to climate fluctuations, detecting and attributing phenological change at broader spatial scales, which can enable a more holistic view of biodiversity change [doi2017macroecological], remains challenging due to widespread data limitations. Some macrophenological studies have managed to overcome these data limitations by using computer vision approaches to extract phenological information from a variety of publicly available data (e.g., digitized herbarium specimens) [williamson2025long]. Yet, most macrophenological studies focus on relatively simple metrics to quantify phenological shifts, such as mean flowering date and flowering duration, as limited intra-annual scale data often precludes more detailed assessments of shift in phenological stages like emergence, flowering onset, seeding, and senescence.
Now however, the wealth of photo-based biodiversity observations produced via community science platforms like iNaturalist can be leveraged to overcome these data shortfalls, if we can extract phenological information from observations in a systematic and reliable manner. Here we test whether INQUIRE-Search can be leveraged to piece together the full life history of an ecologically significant species, common milkweed (Asclepias syriaca), in a specific region (Southern Quebec). We show that INQUIRE-Search can successfully identify observations that cluster into specific phenological stages, emergence, flowering, seeding, and senescence (Figure˜9).
2.4.2 Methodology
We used INQUIRE-Search to retrieve images of target species across multiple life stages:
  1. 1.
    We queried INQUIRE-Search with specific queries for each phenological stage: (1) emergence (“Milkweed germinating or emerging”), (2) flowering (“Milkweed flowering”), (3) seeding (“Milkweed producing seeds or milkweed with seeds”), and (4) senescence (“Milkweed dying or withering or senescence”).
  2. 2.
    We used both a species filter (“Asclepias syriaca”) and a geographic filter (Latitude: 45.03 to 46.54 degrees, Longitude: -74.68 to -71.66 degrees) to isolate observations of common milkweed in southern Quebec.
  3. 3.
    We then used a set of criteria to identify observations that corresponded to specific phenological stages. For emergence, we considered observations with milkweed plants containing less than four pairs of adult leaves. For flowering, we considered observations with open petals. For seeding, we considered observations with open seed pods and visible seeds. For senescence, we considered observations with milkweed plants that no longer exhibited green leaves. For each stage, we queried 200 observations using INQUIRE-Search and marked the observations which corresponded to the matching phenological stage.
Figure 10: (Top) Query & Verification. Top INQUIRE-Search returns for "flowering" phenological stage with incorrect or irrelevant images removed through expert review. (Bottom) Phenological stages occur at significant times throughout the year. (left) density of observations from INQUIRE-Search corresponding to different stages. (right) distribution of observations from INQUIRE-Search in different phenological stages visualized with results from ANOVA/Tukey test.
2.4.3 Results and Analysis
Our analysis shows significant differences in timing for each phenological stage resolved using INQUIRE-Search for common milkweed in southern Quebec. The phenological stage with the best return rate (number of observations matching the queried stage) was flowering (84.5%) followed by seeding (80.5%), senescence (26%), and finally emergence (22.5%). An ANOVA revealed significant differences in stages for the day of year of returned observations, and pairwise differences between stages was tested using a Tukey test. All stages were significantly (p<0.05) different from one another with the exception of emergence and flowering, which showed considerable overlap in timing.
2.4.4 Findings
This case study shows that INQUIRE-Search can be used systematically to resolve detailed phenological information at regional scales for specific species. Some phenological stages (flowering and seeding) were more successfully resolved than others (emergence and senescence). While this could be a limitation of the models ability to identify relevant observations to those stages (the difference between an emerging plant and a young plant is minor) it is also likely due to bias in the community science observations queried, as observations of flowers and seeds are likely more common as they can aid taxonomic identification [iNaturalist202xHighQualityObservations]. Resolving patterns of senescence will also require specific considerations to account for the fact that senescent plants can also be observed long after senescence, evidenced by the fact we identified observations of senescent milkweed throughout the year. Nonetheless, our findings point to the potential utility of INQUIRE-Search to benefit our ability to resolve detailed phenological stages. Considering this sort of fine0scale phenological information would have traditionally required daily or weekly surveys to resolve at the level of a single site, let alone a region like Southern Quebec, our results demonstrate that INQUIRE-Search has the potential to drastically increase our ability to extract phenological information from the wealth of rapidly growing community science observations.
2.5 Whale Re-Identification
2.5.1 Task Description
Animal re-identification (re-ID)—the ability to recognize individual animals across space and time—is central to estimating population size and survival, reconstructing migratory routes, evaluating site fidelity, and monitoring long-term shifts in habitat use [krebs1989ecological]. For wide-ranging marine mammals such as humpback whales, these insights depend on assembling longitudinal photo-ID datasets that link repeated sightings of the same individual across years and ocean locations [katona1981identifying, calambokidis2001movements, martin1984migration]. The pigmentation and trailing-edge patterns on the ventral surface of the tail fluke are often used for identification, and large curated datasets like HappyWhale [howard2018humpback] have been built around these features.
Figure 11: (Left) Re-Identification of humpback whales. Unique Re-Identification of individuals is possible through careful examination of the fluke. (Right) iNaturalist humpback whale images with "fluke" description/tag filtering. Recent iNaturalist humpback whale images, even with the fluke filtering, are not identifiable due to the photographed angle or image quality.
However, scaling re-ID efforts require two components: (1) locating suitable images that clearly show the required body region, and (2) matching those images to known individuals using automated or semi-automated algorithms. While considerable progress has been made on the latter components, with deep re-ID models having shown strong potential for automating fluke matching, locating identifiable images remains a major bottleneck [otarashvili2024multispecies]. Although many fluke images exist in open platforms such as iNaturalist, they are seldom used because there is no effective way to retrieve only those showing a clear, properly oriented ventral fluke suitable for re-identification, while filtering out low-quality or improperly positioned images (see Figure˜11). These identifiable fluke images are scattered across various sources, from community science platforms to tourism photography, each with varying quality, pose, lighting, and availability of metadata.
INQUIRE-Search addresses this gap by enabling targeted retrieval of individually-identifiable images from large, unstructured image collections. Using queries such as "white underside of humpback whale fluke", species filters, and relevancy ranking, the system can surface high-quality fluke images from a global pool of observations. This creates the possibility of augmenting curated catalogs with previously inaccessible sightings, expanding geographic and temporal coverage, and supporting dynamic population monitoring across ocean basins. In this case study, we assess whether INQUIRE-Search can (1) retrieve identifiable humpback whale fluke images from iNaturalist and (2) match them to individuals in the HappyWhale dataset.
2.5.2 Methodology
Figure 12: (Top) Query & Verification. Top INQUIRE-Search returns for "white underside of humpback whale fluke" with incorrect or irrelevant images removed through expert review. Images were marked correct when the right side of the fluke was clearly visible and of sufficient quality for pattern matching. (Bottom) Matching individuals to the HappyWhale database and analysis. A deep learning Re-ID model retrieves candidate matches in the HappyWhale database, which we manually verify, and we then map repeated sightings of confirmed individuals. INQUIRE-Search efficiently surfaces identifiable fluke images from large citizen-science repositories, expanding sighting histories and supporting analyses of movement patterns and habitat use.
We design an approach to efficiently extract identifiable candidate humpback images from the iNaturalist repository. Example retrievals and verifications can be seen at the top of Figure˜12.
  1. 1.
    We queried INQUIRE-Search with a search query including the species name and the body parts required for identification: “white underside of humpback whale fluke”.
  2. 2.
    We analyzed the top 200 images for the search.
  3. 3.
    We manually selected images that provided an unobstructed, clearly visible view of the underside of each whale fluke and were of sufficient quality for analysis. This filtering resulted in 153 images. Most of the 47 excluded images either depicted the wrong part of the whale (e.g., the pectoral fin) or were low quality (blurry or partially obscured).
2.5.3 Results and Analysis
For analysis, we first assess whether any iNaturalist retrievals correspond to known individuals in the HappyWhale dataset. To speed up the matching, we crop each iNaturalist image with Grounding DINO [liu2024grounding] and embed both datasets using a multi-species re-identification model [otarashvili2024multispecies]. For each of the 153 INQUIRE-Search images, we retrieve the top three closest HappyWhale candidates by embedding similarity, then manually review these candidates to confirm matches. Although not conducted by experts, humpback flukes have distinct, easily recognizable patterns that allow reliable identification by non-experts.
Of the 153 candidate images retrieved and verified through INQUIRE-Search, 57 were successfully matched to individuals in the public HappyWhale dataset. Despite inspecting only the top 200 ranked search results, these matches identified 34 unique humpback whales, including multiple sightings of a few individuals, as shown in Figure˜12. These additional sightings extend spatiotemporal coverage of known individuals, demonstrating that INQUIRE-Search can uncover valuable, previously untapped data embedded within large community science repositories.
Figure˜12, which visualizes matched iNaturalist and HappyWhale observations, highlights how this linking expands our ability to track individuals across datasets. Since iNaturalist observations include geolocation metadata, cross-referencing matches provides new spatial information for individuals that were previously represented only in curated catalogs. This capability enables researchers to map additional movement paths, identify unusual sightings, and potentially detect shifts in habitat use when opportunistic images align with structured monitoring efforts.
This proof-of-concept also indicates substantial headroom: evaluating more than the top 200 retrieved images would likely reveal additional high-quality fluke photographs, suggesting that INQUIRE-Search could meaningfully augment re-ID datasets at scale. Importantly, INQUIRE-Search succeeded at retrieving correctly framed, identifiable fluke images despite substantial variation in pose, angle, distance, and environmental settings, which may be conditions under which conventional metadata searches would fail.
2.5.4 Findings
While small in scale, this case study demonstrates that images can be surfaced by INQUIRE-Search, then linked to known individuals using re-ID, providing valuable new information for population monitoring. INQUIRE shows strong potential for retrieving high-quality, correctly posed images that capture identifiable features, which enables the integration of valuable data from previously disconnected sources. However, its broader utility is constrained by the identification step, which still depends on a pre-existing labeled dataset and a trained model for matching. Because INQUIRE cannot assign individual identities on its own, its use is more limited when building datasets from scratch or incorporating previously unseen individuals. In addition, retrieval performance reflects the spatial and behavioral biases of community-sourced photographs, meaning the tool will preferentially surface observations from well-sampled areas and charismatic encounters rather than providing an unbiased representation of whale distribution or behavior.
3 Discussion
Across our five case studies, we show that INQUIRE-Search accelerates ecological discovery by transforming community-sourced imagery into analyzable evidence. The system enables rapid, expert-guided exploration of ecological questions that would otherwise require extensive field effort or months of manual screening. In each application, INQUIRE-Search helped resolve major data gaps by efficiently recovering information at scale, either through supporting the study of spatiotemporal patterns (bird diet, post-fire regrowth, and phenology), uncovering rare or hard-to-detect events (wildlife mortality), or identifying observations of sufficient quality for downstream analysis (whale re-identification). Together, these examples show how INQUIRE-Search can be used to not only accelerate data discovery but also to fill critical temporal, spatial, and observational gaps across ecological domains.
3.1 Rethinking the Scientific Process in the Age of Image Search
With the interactive, iterative approach to data discovery enabled by INQUIRE-Search, scientists remain the drivers of inquiry, interpretation, and verification. Incorporating such a tool with an ecological research workflow requires rethinking experimental design. The following discussion highlights alternate parameters to "tune" in the scientific process, from identifying data-driven problem settings with the appropriate scope and iteratively crafting robust queries, to capping verification effort and quantifying uncertainty.
3.1.1 Which questions are well-posed for INQUIRE-Search?
Determining which scientific questions are best-posed to be addressed with INQUIRE-Search becomes essential. The tool is fundamentally limited by the images that exist in the underlying repository. Questions targeting extremely rare events, under-observed species, or obscure behaviors may yield insufficient data for robust analysis. For example, in the avian diet case study, species with few iNaturalist records produced too few relevant images of feeding behaviors to support quantitative comparisons across time. Conversely, when data are overly-abundant, as in the wildlife mortality study, search still improves efficiency but does not eliminate the need for substantial verification. Additionally, a query is only valuable if the target feature can be unambiguously determined from an image. Prompts such as "robin carrying a worm" or "milkweed with seed pods" are verifiable because the relevant features are visibly explicit. In contrast, prompts such as "bird preparing to migrate" or "tree under stress" are too abstract to evaluate consistently from a still image. Well-posed questions therefore target phenomena that are (1) sufficiently represented in the data pool, (2) not so broad or ubiquitous that they produce large volumes of ambiguous or redundant results, and (3) visually verifiable.
3.1.2 How to refine a question: iterating on prompts
Search-based discovery depends on clear, specific prompts. Precise language reduces ambiguity in the embedding space and improves the relevance of returned images. Specificity also increases the the verifiability of the outputs: "robin pulling a worm from the soil" will yield cleaner results than "bird foraging," which may return heterogeneous behaviors. Effective prompt-crafting often requires iteration, especially when early results are too broad (refine behavior, context, or anatomical cues) or results contain few matches (broaden the phrasing or adjust the filters). This emphasizes the importance of developing appropriate query design and iterative refinement when using AI-assisted systems like INQUIRE-Search, analogous to testing refining sampling strategies in field studies.
3.1.3 Managing Effort, Ranking, and Verification
This workflow introduces a new design choice: in addition to formulating effective queries, researchers must also decide how far down the ranked search results to inspect. In this way, "effort" becomes tunable. Reviewing only the top-ranked subset often recovers the most relevant observations with minimal redundancy, as shown in the dietary analysis case study, where experts verified only first 200 images recovered to obtain a sufficient quantity of data.
3.1.4 Identifying Uncertainty and Bias with INQUIRE-Search
INQUIRE-Search inherits the limitations present in the pool of data it retrieves from. Community science platforms such as iNaturalist contain strong observer biases: charismatic or unusual events are overrepresented, while cryptic interactions or remote habitats are under-documented [dimson2023observer]. The accessibility of a study site/subject influences how detectable it is—urban and roadside species are recorded far more often than taxa in remote forests, offshore waters, or high-latitude environments [dimson2023observer]. Methods for observer-effort corrections exist to quantify and mitigate these biases [robinson2018correcting, guilbaultexplorers, mondain2024adaptive, padilla2024assessing], but their effectiveness at a global scale remains limited because they require structured sampling, repeated surveys, or explicit effort measurements that community science data often lack.
Model-driven biases add a second layer of uncertainty. Some prompts may be systematically harder for the embedding model to resolve, such as those that involve examples with low contrast, occlusion, rare species, and unfamiliar backgrounds, introducing an additional layer of uncertainty. VLMs also struggle with prompts requiring complex relational or compositional understanding [alhamoud2025visionlanguagemodelsunderstandnegation, thrush2022winogroundprobingvisionlanguage]. Moreover, recent work shows that VLMs are both miscalibrated (where the expressed confidence does not match true accuracy) [guo2017calibration] and biased (the model defaults to memorized training-set associations instead of analyzing visual evidence) [vo2025vision]. Adapting existing statistical tools for bias estimation and correction [narduzzi2014inverse, geifman2017selective, karimi2023quantifying] could provide a foundation for quantifying uncertainty in discovery-driven search.
3.2 Expanding Beyond iNaturalist
Although our case studies drew on iNaturalist for its scale and ecological breadth, the open-source design of the INQUIRE-Search codebase makes it adaptable to any large image collection. Early qualitative experiments suggest that the system has potential for other ecological image repositories, such as camera trap datasets. We see INQUIRE-Search less as a single-platform tool and more as a general methodology to enable scientists to get more from their data, flexibly and quickly. Additionally, because the system is open-source and resource-efficient, it has the potential to be deployed widely, enabling scientists, NGOs, and local communities to harness image search for ecological monitoring and hypothesis generation on their own databases.
4 Conclusion
Growing ecological image databases are a barely-tapped scientific resource, compared to the hidden knowledge they hold. We show that our carefully designed and engineered system, INQUIRE-Search, can enable scientists to work iteratively via open-ended natural language search and verification to rapidly and scalably discover diverse secondary ecological signals, from bird diet to plant phenology. These capabilities point toward a new paradigm of experiment design and data collection in ecology, where hypothesis generation, testing, and data curation occur within a unified, interactive search system over already-collected data. INQUIRE-Search reframes community-sourced imagery from incidental records of species occurrences into an accessible and systematic source of ecological evidence. This perspective democratizes access to data-driven discovery while also raising new challenges around reproducibility, rigor, and inference. It calls on ecologists to integrate search tools into study design: deciding what constitutes a good query, how to measure uncertainty, and how to validate findings across scales. When integrated with careful experimental design, INQUIRE-Search complements traditional ecological methods, offering a scalable path to accelerate discovery, expand participation, and deepen our understanding of biodiversity in a rapidly changing world.
5 Appendix
5.1 System Architecture and Design Motivation
INQUIRE-Search is a tool designed to enable scientists to quickly and easily discover data from within a large ecological database such as iNaturalist. This challenge of surfacing data of interest from a large pool is commonly referred to as "image retrieval" in the AI community, and INQUIRE-Search specifically addresses two fundamental image retrieval challenges to ensure the system is high-performing in terms of both quality and speed. First, INQUIRE-Search enables scientists to search for data using open-ended text, which significantly expands the searchable possibilities beyond existing metadata within the database. Second, INQUIRE-Search is engineered to power this search over very large databases efficiently, without requiring significant computational resources. This efficiency is a key feature of the design, letting scientists iterate on their searches with the system in real-time to discover their data of interest. The system architecture of INQUIRE-Search reflects both of these priorities, combining state-of-the-art vision-language models with efficient indexing and memory management techniques to deliver responsive searches across hundreds of millions of images on modest hardware (4 vCPUs with 32GB RAM, roughly the capacity of a single laptop).
5.2 Search Index
Searching for relevant data using text is accomplished by first embedding the text query using the same SigLIP model used to embed the iNaturalist images. We then calculate the vector similarity between that new text embedding and each image embedding and use that similarity score to rank the full set of images to surface possible images of interest. With a smaller database, this entire process can be accomplished directly within the memory of a modest computer, but when scaling to the hundreds of millions of images in iNaturalist the set of image embeddings becomes large ($\sim$200GB), requiring a different approach to quickly find the most similar images to a new text query.
Vector databases, which use a specialized indexing system that allows fast search over millions of high-dimensional vectors, [decastro2018detecting, singla2021experimental, fang2017topology] are specifically designed to facilitate efficient similarity-based search. We utilize FAISS [douze2025faiss] to create an approximate nearest neighbor index that is small and fast to query. Our FAISS index is generated and tuned using the Autofaiss library, a widely-used library for efficient nearest-neighbor search [webster2023duplication]. The final index is memory-mapped (stored on disk but accessed as if it were in memory, enabling fast search with limited RAM), uses 36GB of storage, and enables sub-500ms search times while maintaining high retrieval accuracy.
In addition to the visual embedding index, our search tool supports several metadata-based mechanisms to filter by:
  1. 1.
    Taxonomy: We map each taxon (species, genus, family, etc.) to a list of image IDs known to contain that entity, using taxon labels provided for each image from iNaturalist. This allows for efficient subset selection when users apply taxonomic filters to their queries.
  2. 2.
    Temporal Range: Users can filter search queries by month.
  3. 3.
    Geographic Extent: Using approximate image location data, optionally provided by iNaturalist contributors, users can filter images to a specific geographic area using latitude-longitude bounds.
5.3 Image and Text Embeddings
The core search capability behind INQUIRE-Search is powered by such a vision-language model. Not all vision-language models are equally capable of ranking ecological data, particularly for scientific queries. We leveraged our previous work developing the INQUIRE-Benchmark, which specifically evaluates models for ecological image retrieval tasks, to select SigLIP-So400m-384-14 [zhai2023sigmoid] as our embedding model, due to both its strong performance on the benchmark and its reasonable tradeoff between speed and accuracy; other models like ViT-H-14 [radford2021learning] are slightly more accurate at retrieving relevant data but are significantly larger, and thus slower. We processed 300M images sourced from iNaturalist through the visual encoder component of the selected SigLIP model to build the backbone of the INQUIRE-Search tool.
5.4 Search Interface
INQUIRE-Search provides a web-based interface (Fig. 13) that surfaces the search index and filters to scientists. The interface includes a query field for natural language queries (e.g., "California condor with a green ‘26’ tag on its wing") and controls for taxonomic, temporal, or geospatial filters. The interface displays a grid of thumbnails of the images returned from the user-provided query, ordered by relevance to the query. The images are clickable, allowing the user to mark images for further analysis. An expanded view of a selected image shows a full-resolution image with complete metadata.
Figure 13: INQUIRE-Search user interface. INQUIRE-Search provides an user-friendly and intuitive interface that allows ecologists to query hundreds of millions of natural-world images. This example illustrates a query for "American robin eating," with species and temporal filters applied, enabling rapid retrieval, inspection, and verification of relevant observations.
5.5 Data Export
When the user exports the data, a CSV file is prepared for download. This file contains a row for each image which includes a field for whether an image was marked by the user, detailed metadata, and a link to the associated iNaturalist observation.
5.6 Data Analysis
Once verified, these retrieved observations form the foundation for a wide range of statistical analyses, demonstrated in the following section. This workflow allows experts to effectively "design" their search strategy and specify what ecological signals to target, while identifying or mitigating relevant sources of uncertainty and bias.