### v3

# INQUIRE-Search: A Framework for Interactive Discovery in Large-Scale Biodiversity Databases

Abstract
  1. 1.
    Community science platforms such as iNaturalist contain hundreds of millions of biodiversity images that capture ecological context on behaviors, interactions, phenology, and habitat. Yet ecological workflows to surface these data rely on metadata filtering or manual inspection, leaving this "secondary information" largely inaccessible at scale.
  2. 2.
    We introduce INQUIRE-Search, an open-source system that uses natural language to enable scientists to rapidly search within an ecological image database for specific concepts, verify and export relevant observations, and use these outputs for downstream scientific analysis. Compared to traditional methods, INQUIRE-Search takes a fraction of the time, opening up new possibilities for scientific questions that can be explored.
  3. 3.
    Through five case studies, we demonstrate the range of applications INQUIRE-Search can support, from seasonal variation in behavior across species to forest regrowth after wildfires. These examples demonstrate a new paradigm for interactive, efficient, and scalable scientific discovery that can begin to unlock previously inaccessible scientific value in large-scale biodiversity datasets.
  4. 4.
    Finally, we highlight how AI-enabled discovery tools for science require reframing aspects of the scientific process, including experiment design, data collection, survey effort, and uncertainty analysis.
Keywords community science, computer vision, deep learning, natural language queries, secondary data, vector search
1 Introduction
1.1 The Need for Tools to Make Ecological Databases Discoverable
Image data have become a key component of ecological research, collected to document and analyze biodiversity, species interactions, and environmental changes across spatial and temporal scales [oliver2023camera, kays2020born]. With the widespread use of digital photography and automated sensors, community science platforms like iNaturalist [iNaturalist2024] now contain hundreds of millions of image-based species observations [callaghan2021three].
Currently, the scientific use of such databases has primarily focused on species occurrences [kumar2019using, humphreys2019seasonal], but many ecological questions, such as biotic and abiotic interactions, resource use, and habitat associations, require information that extends beyond species identification and location alone. We refer to this contextual information as secondary data: ecologically relevant features that extend beyond the primary observation [pernat2024overcoming, marques2024retrieving, davison2025automated]. Images frequently capture organisms in their microhabitats, with co-occurring species, observable behaviors, or signs of disturbance.
[…]
Figure 1: INQUIRE-Search pipeline overview. Users begin with a scientific question and iteratively (1) query and prioritize relevant iNaturalist images, (2) refine and verify results with expert inspection, and (3) export the curated data for downstream analysis. The process supports returning to earlier stages as needed until sufficient data are collected.
1.2 Utilizing Machine Learning to Process Large Scale Data
Recent advances in computer vision provide valuable tools for dataset curation and filtering, and automated species detection models are widely used in ecology [norouzzadeh2018automatically, willi2019identifying]. However, supervised learning approaches to train these models require large, well-labeled datasets, which are often hard to acquire for ecological information of interest. This closed-set design limits the ability of supervised models to generalize beyond fixed categories and constrains their use in open-ended ecological exploration.
Vision-language models (VLMs) offer a powerful alternative. Trained on large image-text datasets, these models map images and language into a shared high-dimensional representation space where semantically similar concepts lie close together [radford2021learning]. For example, an image of a hermit crab wearing plastic as a shell would be embedded into this space closer to the corresponding text "a hermit crab with plastic trash as a shell," and far from unrelated descriptions such as "a yellow car," as seen in Fig 2. VLMs therefore enable us to search for relevant images in a database through a simple similarity search—ranking each image in the database based on its distance to the embedding of a text query.
Figure 2: (Top) Joint vision-language representation space. VLMs are trained to learn a joint embedding space between vision and text modalities, where similar semantic concepts are stored closer together. (Bottom) Image-text search. At search-time, the input text query is embedded by the text encoder of the VLM. Then the query embedding vector is compared against the pre-computed iNaturalist image embeddings to retrieve the most relevant images.
Embedding-enabled search with VLMs has already shown promise for ecological applications. WildCLIP [gabeff2024wildclip] fine-tuned a VLM on camera trap images and ecology-specific language descriptions. From training on large data, this model was able to generalize from concepts seen during training to search large wildlife datasets for new behaviors or habitat details. The INQUIRE Benchmark [vendrow2024inquire] evaluates state-of-the-art VLMs on expert-level open-ended ecological queries across five million iNaturalist 2024 images. Both works clearly demonstrate the potential of such an approach to unlocking new knowledge in databases. Realizing this potential requires a system that connects the broad capabilities of VLMs with the practical needs of ecological research, enabling scientists to make effective use of large image repositories.
To address the need, we introduce INQUIRE-Search, a system for efficient, expert-driven retrieval from large ecological image datasets. As illustrated in Fig 1, the system facilitates an iterative cycle in which researchers pose questions in plain language, retrieve semantically relevant images, assess and refine those results based on ecological validity, and ultimately export curated observations for downstream analysis.
2 Materials and Methods
2.1 INQUIRE-Search System Architecture
INQUIRE-Search is an open-source system for interactive, expert-driven retrieval from large ecological image datasets. It integrates a vision–language embedding model, a high-performance vector index, and a lightweight, browser-based interface to support efficient natural-language search, verification, and data export.
The system uses a state-of-the-art Vision–Language Model (VLM), SigLIP-So400m-384-14 [zhai2023sigmoid], to embed both images and natural-language queries into a shared high-dimensional semantic space. All iNaturalist [iNaturalist2024] images were preprocessed to a uniform format and embedded using the SigLIP image encoder. These embeddings were stored in a FAISS (Facebook AI Similarity Search) [douze2025faiss] index, enabling sub-second similarity search across approximately 300 million images.
At search time, a user-provided text query is embedded and used to rank images in the FAISS index by cosine similarity. The browser-based interface supports metadata filtering (taxonomy, geography, and date), rapid inspection and expert verification of retrieved images, and export of curated observations with complete iNaturalist metadata. Verified image sets can be exported as CSV files, including observation IDs, coordinates, timestamps, taxonomic information, and file URLs, for downstream analysis. Full implementation details are provided in the Appendix.
2.2 INQUIRE-Search Pipeline Overview
We follow a standardized human-in-the-loop workflow using INQUIRE-Search.
  1. 1.
    Query and Prioritize: We composed natural-language queries tailored to each ecological question and applied optional metadata filters (species, date, location). All prompts are listed in Table 1, with complete information on the filters, number of retrievals inspected, and number of samples marked for analysis provided in the Appendix.
  2. 2.
    Retrieve and Verify: For each query, the system returned top-ranked images. Depending on the case study, we screened between 200 and 500 top retrievals for relevant data. A reviewer examined each image and labeled it informative only if it met the case-specific criteria.
  3. 3.
    Export and Analyze: Informative images and their metadata were exported as CSV files for downstream analysis. Analytical procedures were specific to each case study and are described in Section 2.3.
Case study INQUIRE-Search Prompt
Seasonal variation in bird diets Prompts of the form "<species> with <diet type> in its mouth" using: Species (common names): Gray-cheeked Thrush, Ancient Murrelet, American Tree Sparrow, Red-bellied Woodpecker, American Robin Diet types: invertebrate, vertebrate, seed, fruit, nectar and pollen, carrion or animal decay, other plant matter Total: 5 species × 7 diet types = 35 prompts.
Post-fire forest regrowth "Young coniferous trees in burned forest"
"Young deciduous trees in burned forest".
Wildlife mortality "Dead bird"
Plant Phenology "Milkweed germinating or emerging"
"Milkweed flowering"
"Milkweed producing seeds"
"Milkweed dying or withering / in senescence".
Whale re-identification "White underside of humpback whale fluke".
Table 1: Prompts used for retrieval across the five INQUIRE-Search case studies.
2.3 INQUIRE-Search Case Studies
We collaborated with a diverse set of ecological experts to define a series of scientific questions that would potentially be addressable with information hidden within the iNaturalist database. Together, we used INQUIRE-Search to address each of these questions, designing each experiment to test both the usability and utility of the tool and test the limits of what insights the system could help uncover. Each case study explores a distinct ecological question: (1) seasonal variation in bird diets, (2) post-fire forest regeneration, (3) spatio-temporal patterns of wildlife mortality, (4) plant phenology across seasonal cycles, and (5) individual Re-ID in humpback whales. These examples show how natural language-guided image search and analysis can provide a scalable and flexible approach for extracting ecological insights from large, unstructured image datasets.
2.3.1 CS1: Seasonal Variation in Bird Diets
Overview Diet is a fundamental response and effect trait that shapes vertebrate survival, fitness, trophic position, and ecological interactions, and it varies widely across species and environmental conditions [burin2016omnivory, sibly2012energetics, belmaker2012global]. Although large dietary databases exist [wilman2014eltontraits], seasonal diet information remains sparse: in SAviTraits 1.0 [murphy2023SAviTraits], only about 10% of more than 10,000 bird species are recorded as exhibiting seasonal dietary shifts, a pattern likely driven by data gaps rather than true biological invariance. Extracting such information from community-science imagery is also nontrivial: among over 23,000 relevant iNaturalist observations of American Robin, only 44 are tagged “feeding,” and none specify the type of food consumed. This case study evaluates whether INQUIRE-Search can systematically recover feeding events across seasons from community-sourced photographs and reproduce known dietary patterns. We focus on five species – Gray-cheeked Thrush (Catharus minimus), Ancient Murrelet (Synthliboramphus antiquus), American Tree Sparrow (Spizelloides arborea), Red-bellied Woodpecker (Melanerpes carolinus), and American Robin (Turdus migratorius – which have high dietary certainty scores in SAviTraits [murphy2023SAviTraits], providing a robust reference for comparison.
Search Strategy: For each target species, we construct natural language search queries specifying the bird species and diet type (e.g., "American robin with invertebrate in its mouth"), using the dietary classification system used in SAviTraits. This classification system includes seven dietary categories: (1) invertebrate, (2) vertebrate, (3) seed, (4) fruit, (5) nectar and pollen, (6) carrion or animal decay, and (7) other plant matter. Each diet category was queried separately for summer (June–August) and winter (December–February). Taxonomic filters restricted results to the target species.
Verification: An image was considered informative only when a clearly identifiable food item was visible in the bird’s bill. Top 500 images were screened by a reviewer for every species-diet type-season combination.
Analysis: For each combination, we counted the marked images and calculated the proportion of informative images belonging to each diet category. These proportions were then compared qualitatively with published seasonal diet compositions from SAviTraits.
2.3.2 CS2: Post-fire Forest Regrowth
Figure 3: (Left) Forest regrowth following the 2012 High Park Fire in Colorado. With wildfires becoming more frequent and severe, fine-scale characterization of vegetation recovery is critical for understanding forest resilience. (Right) MTBS burn severity categories. We aggregate the burn categories from 30m Landsat resolution to a spatial resolution matching iNaturalist metadata (0.01, $\sim$1100m).
Overview: Understanding how forests recover following fire is essential for assessing long-term ecosystem resilience, particularly in the western United States where wildfire has been the dominant disturbance for millennia [whitlock2008long, marlon2012long] and many forests are fire-adapted [keeley2011fire]. Recent warming and drying trends associated with climate change, coupled with the legacy of fire suppression, are driving fires that are more frequent, severe, and intense [abatzoglou2016impact, juang2022rapid], exceeding the adaptive capacity of many forest systems by limiting post-fire regeneration and altering species composition [harvey2016burn, davis2023reduced]. Despite the ecological importance of post-fire recovery for understanding forest resilience, successional trajectories, and long-term ecosystem change, existing approaches provide limited capacity to assess fine-scale regrowth: satellite data often lack the spatial or spectral resolution needed to detect young trees or differentiate species [xu2021spatial, kiel2022trees], while field surveys are constrained by cost and limited spatial coverage. Because community-sourced photographs can reveal early regeneration that is invisible to both satellites and sparse field campaigns, this case study evaluates whether INQUIRE-Search can identify young coniferous and deciduous individuals within the 2012 High Park Fire perimeter and examine how their occurrence varies across burn severity classes (Fig. 3).
Search Strategy: We targeted two functional groups (young coniferous trees and young deciduous trees), using text queries such as "young coniferous trees in burned forest" and "young deciduous trees in burned forest". Spatial filters restricted observations to the High Park Fire perimeter: latitude (40.57 to 40.75 N) and longitude (105.18 to 105.54 W) (Fig 3). Temporal filters restricted images to post-fire dates. After retrieval, we applied a finer spatial filter using reported image coordinates, retaining only images within the High Park MTBS fire perimeter via the sf package in R3 [pebesma2018simple].
Verification: An image was marked informative when it contained the forest community of interest (i.e., an individual of the target forest type could be clearly distinguished). We allowed individual images to be marked as informative in multiple searches if the image contained individuals representing both forest types (e.g. a coniferous and deciduous tree seedling). While processing images for searches related to trees, we also decided to include trees and shrubs within the same target forest type as it was difficult to distinguish young trees from shrubs in images. Top 200 images were inspected.
Analysis: Coordinates of informative images were combined with MTBS burn severity categories by aggregating the 30 m burn data to the 0.01 precision typical of iNaturalist coordinates and assigning each location the modal severity class using the terra package in R4 [hijmans2022package]. We then counted marked images by forest type and burn severity to evaluate how regeneration outcomes varied across burn categories.
2.3.3 CS3: Wildlife Mortality
Figure 4: (Left) Wildlife mortality trends vary by season and location. Globally, collisions with human-made structures such as windows are a major cause of bird deaths. (Right) Comparing avian mortality in rural vs urban regions. Using INQUIRE-Search, we compared image-based evidence of avian mortality between urban (Boston, MA) and rural (Pioneer Valley, MA) sites to examine how seasonal risks differ across anthropogenic contexts.
Overview: Understanding patterns of avian mortality is fundamental to conservation efforts, particularly given the loss of nearly three billion birds in North America since 1970 [rosenberg2019decline]. Mortality risk varies seasonally as birds encounter different threats across the annual cycle [loss2015direct, marra2015call], and collisions with human-made structures–one of the leading anthropogenic causes of bird death [loss2014estimation]–show strong seasonal peaks during migration [riding2021multi, scott2023causes] and vary across landscapes due to differing local hazards and bird abundances [hager2017continent]. Yet documenting spatiotemporal patterns of mortality is challenging because mortality events are rare, and alternative monitoring methods such as marking or tracking individuals are resource-intensive and limited in scale [yanco2025tracking]. Since large-scale image repositories could reveal where and when birds die, this case study uses INQUIRE-Search to investigate seasonal mortality dynamics in an urban (Boston, MA) and rural (Pioneer Valley, MA) setting and compare relative mortality between these contrasting anthropogenic contexts (Fig 4).
Search Strategy: We queried for "dead bird" filtering by each of the two equal-sized bounding boxes using location filters. We use latitude (42.31 to 42.38 N) and longitude (-71.14 to -71.01 W) for the urban study area, and latitude (42.31 to 42.38 N) and longitude (-72.64 to -72.51 W) for the rural study area.
Verification: We excluded images of live birds, single features, non-avian taxa, and duplicate observations of the same event (duplicate sharing species, month, latitude and longitude) to identify true avian mortality events.
Analysis: To control for differing bird abundance, observation effort, and mortality among sites, we calculate the mortality index defined as:
$\text{MortalityIndex}_{m,s}=\log_{2}\!\left(\frac{R_{m,s}}{\overline{R}_{s}}\right),$
where the monthly mortality rate is
$R_{m,s}=\frac{D_{m,s}}{O_{m,s}},$
and the mean monthly mortality rate for site $s$ is
$\overline{R}_{s}=\frac{1}{12}\sum_{m=1}^{12}R_{m,s}.$
Here, $D_{m,s}$ is the mortality count and $O_{m,s}$ is the observation count for month $m$ at site $s$. Values above zero indicate fold increases in mortality relative to the site’s annual mean, while values below zero indicate fold decreases. The total observation count was extracted using the rinat package in R [barve2022rinat].
2.3.4 CS4: Resolving Plant Phenology
Overview: Studying the timing of life-history events is a primary way researchers detect biodiversity responses to global change [parmesan2003globally], as phenological shifts in processes such as germination, flowering, and senescence can serve as early warning signals that precede demographic, genetic, or geographic change. While fine-scale phenological datasets collected through repeated field surveys provide valuable local insight [austin2024climate], detecting and attributing phenological change at broader spatial scales remains challenging due to limited temporal resolution and data availability [doi2017macroecological]. Although some macrophenological studies have used computer vision to extract phenological information from digitized specimens [williamson2025long], most rely on coarse metrics such as mean flowering date because intra-annual data on distinct phenophases are sparse. The growing volume of geotagged plant observations on community-science platforms like iNaturalist offers a path to overcoming these limitations, provided phenological stages can be systematically identified from photographs. In this case study, we test whether INQUIRE-Search can recover detailed phenophases—emergence, flowering, seeding, and senescence—for common milkweed (Asclepias syriaca) in southern Québec (Fig 5).
Figure 5: Four phenological stages of common milkweed. Example observations identified using INQUIRE-Search that correspond to the four phenological stages analyzed in this study: emergence, flowering, seeding, and senescence.
Search Strategy: We created stage-specific text queries targeting (1) emergence ("Milkweed germinating or emerging"), (2) flowering ("Milkweed flowering"), (3) seeding ("Milkweed producing seeds or milkweed with seeds"), and (4) senescence ("Milkweed dying or withering or senescence"). We also used a species filter ("Asclepias syriaca") and a geographic filters for latitude (45.03 to 46.54 N) and longitude (-74.68 to -71.66 W) to limit observations to southern Quebec.
Verification: Top 200 images were labeled as informative according to strict morphological criteria:
  • Emergence: <4<4 pairs of adult leaves
  • Flowering: presence of open petals
  • Seeding: open pods with visible seeds
  • Senescence: absence of green leaves
Analysis: Observation dates were converted to Day-of-Year (DOY). Mean DOY values were compared among stages using ANOVA, followed by Tukey’s HSD to identify pairwise differences.
2.3.5 CS5: Whale Re-Identification
Overview: Animal re-identification – the ability to recognize individual animals across space and time – is essential for estimating population size and survival, reconstructing migratory routes, evaluating site fidelity, and monitoring long-term shifts in habitat use across different species and ecosystems [krebs1989ecological]. For wide-ranging marine mammals such as humpback whales, these insights depend on longitudinal photo-ID datasets linking repeated sightings across years and ocean basins [katona1981identifying, calambokidis2001movements, martin1984migration], typically using pigmentation and trailing-edge patterns on the ventral fluke [howard2018humpback]. Although many fluke images exist in open repositories like iNaturalist, scaling re-ID remains constrained by the difficulty of locating images that show a clear, properly oriented fluke suitable for identification, as existing filters cannot reliably distinguish low-quality, mis-angled, or incomplete views (Fig 6) [otarashvili2024multispecies]. Because identifiable fluke photographs are scattered across diverse image sources with variable pose, lighting, and metadata, this case study evaluates whether INQUIRE-Search can surface individually identifiable humpback whale flukes from unstructured community-science imagery and match them to known individuals in the HappyWhale dataset.
Figure 6: (Left) Re-Identification of humpback whales. Unique Re-Identification of individuals is possible through careful examination of the fluke. (Right) iNaturalist humpback whale images with "fluke" description/tag filtering. Recent iNaturalist humpback whale images, even with the fluke filtering, are not identifiable due to the photographed angle or image quality.
Search Strategy: We queried for "white underside of humpback whale fluke" with the humpback whale species filter.
Verification: Images were considered informative only if the ventral fluke was clearly visible, unobstructed, and in focus. Top 200 retrievals were inspected.
Analysis: For analysis, we assess whether any iNaturalist retrievals corresponded to known individuals in the HappyWhale dataset. To accelerate matching, we cropped each image using Grounding DINO [liu2024grounding] and embedded both the iNaturalist and HappyWhale datasets using a multi-species re-identification model [otarashvili2024multispecies]. For each of the verified INQUIRE-Search images, we retrieved the top three closest HappyWhale candidates by embedding similarity and manually reviewed these candidates to confirm matches. Although not conducted by experts, humpback flukes possess distinct and easily recognizable pigmentation and trailing-edge patterns that allow reliable identification by non-experts.
3 Results
3.1 CS1: Seasonal Variation in Bird Diets
Results and Analysis
Our analysis revealed substantial variation in the availability of dietary information across the five study species shown in Fig 7. American Robin yielded the highest number of informative (i.e., containing dietary information) images (669 total), followed by Red-bellied Woodpecker and American Tree Sparrow (395 and 188, respectively), while for both Ancient Murrelet and Gray-cheeked Thrush we found a negligible number of images containing dietary information. Among the species for which dietary information was available, seasonal patterns in image availability were also evident: American Robin had more informative images in summer than winter, while Red-bellied Woodpecker and American Tree Sparrow had more images available for winter months.
Figure 7: (Top) Query & Verification. Filtering top INQUIRE-Search outputs for "Red-bellied Woodpecker" and "Invertebrate." (Bottom) Diet comparisons. INQUIRE-Search yields dietary patterns that align closely with those documented in the SAviTraits reference dataset.
For the three species with sufficient data, dietary composition derived from INQUIRE-Search closely matched reference data from SAviTraits (Fig 7). For example, in summer, American Robin exhibited an even split between fruit and invertebrate consumption (50% each) in both datasets, while Red-bellied Woodpecker showed a diet dominated by seeds (50% in INQUIRE-Search vs. 40% in SAviTraits), with smaller contributions from invertebrates, fruits, and plant material. In winter, all three species exhibited diets strongly dominated by a single food type, and we again observed close agreement between INQUIRE-Search outputs and SAviTraits data.
Findings
We demonstrate that our method can effectively reproduce dietary patterns documented in established literature, provided that sufficient image data are available. The close alignment between INQUIRE-Search-derived avian diets and SAviTraits reference data highlights the potential of community science imagery to yield accurate ecological insights when properly analyzed through a systematic search methodology. However, this approach is most effective for commonly photographed species such as American Robin and Red-bellied Woodpecker. Two of the species we examined lacked sufficient image data for meaningful analysis despite being well-studied in the scientific literature and represented in databases like SAviTraits, underscoring a fundamental constraint of community-sourced data. This limitation is likely even more pronounced for understudied species—precisely those for which new dietary data are most needed—because they are also underrepresented in iNaturalist and thus rarely retrievable with INQUIRE-Search.
3.2 CS2: Post-fire Forest Regrowth
Figure 8: (Top) Query & Verification. Filtering top INQUIRE-Search returns for "young coniferous trees in burned forest." (Bottom) Tree observations across burn severity. Distribution of observed young coniferous and deciduous trees in community science collected images show strong relationships with burn severity regions.
Results and Analysis
Of the 200 images exported from INQUIRE-Search for each search, there were 100 (coniferous) and 112 (deciduous) images with reported coordinates within the High Park burn perimeter. We marked more informative images for deciduous forest species than coniferous (78 vs. 45 images, respectively). We also marked five images depicting both coniferous and deciduous individuals, indicating areas that could recover into mixed forest communities.
Among images marked for analysis, we found strong relationships between burn severity and post-fire regeneration of both coniferous and deciduous forest types. For both forest types, most marked images of regenerating vegetation occurred in areas classified as unburned-to-low or low burn severity, with few observations in moderate or high severity areas (Fig 8, left). However, coniferous regeneration showed a stronger negative association with increasing burn severity: 93.3% of coniferous images were located in unburned-to-low or low severity areas, with only 4.4% and 2.2% in moderate and high burn severity areas, respectively (Fig 8, right). In contrast, deciduous regeneration was somewhat more evenly distributed, with 84.6% of marked images in unburned-to-low or low severity areas, and a higher proportion (7.6% each) in areas burned at moderate and high severity (Fig 8, right). These patterns suggest that post-fire forest regeneration, as captured by community-science images, is more common in lower severity areas, with coniferous species showing greater sensitivity than deciduous species.
Findings
In this case study, we demonstrate that INQUIRE-Search can be used to effectively identify post-fire forest regeneration patterns across burn severity gradients. Our approach successfully distinguished between regenerating coniferous and deciduous forest types and provided spatially explicit insights into post-disturbance recovery within the first decade after the High Park fire. Our findings are consistent with previous research documenting that increased fire severity decreases conifer regeneration likelihood both regionally and specifically after the High Park fire [davis2023reduced, wright2017post].
Several limitations remain. Image coordinates are rounded to 0.01 to match the limiting spatial precision of many iNaturalist records, but this prevents fine-scale matching to burn severity or site attributes such as seed source or soil moisture. Classification from user-submitted photographs also introduces uncertainty, as landscape images often obscure species identification, and deciduous seedlings, saplings, shrubs, and herbaceous plants can appear morphologically similar. Finally, iNaturalist data are opportunistic and biased toward accessible, aesthetically appealing areas, which may result in severely burned sites being underrepresented (Fig 3). Thus, while INQUIRE is a powerful supplemental tool for assessing post-fire recovery, its results should be interpreted alongside field data and remote-sensing derived products to minimize bias.
3.3 CS3: Wildlife Mortality
Figure 9: Seasonal patterns in relative bird mortality for an urban (Boston, blue) and rural (Pioneer Valley, orange) region of Massachusetts. Mortality is shown as a $\log_{2}$ index scaled to each site’s annual mean (0), with units representing fold-change. Positive values indicate above-average mortality, negative values below.
Results and Analysis
The results show distinct patterns between urban and rural settings. Boston exhibits notable peaks in mortality during the spring and fall, moderate mortality rates during the summer, and low mortality rates in winter. In contrast, the Pioneer Valley shows elevated mortality rates in summer and fall, moderate mortality rates in early winter, and low mortality rates in late winter and spring. The high mortality rates at both sites in September and October are consistent with expectations of peak migratory passage—and the concomitant risk of building collisions—being highest in the fall in the eastern United States [horton2019bright]. Interestingly, Boston exhibited higher mortality rates during spring migration which may reflect differential migratory passage rates or other differences in site-specific risk factors [scott2023causes]. Finally, the Pioneer Valley did not exhibit the winter drop in mortality rates that we observed in Boston. While this could reflect differences in the avian communities or risks between the two sites, we cannot rule out the possible role of increased detection probability due to reduced vegetative cover disproportionately affecting the more rural Pioneer Valley.
Findings
We demonstrate that INQUIRE-Search can detect distinct spatiotemporal patterns in avian mortality through analysis of community-sourced photographs. The contrasting seasonal dynamics between urban (Boston) and rural (Pioneer Valley) sites provide a compelling demonstration of the importance of anthropogenic context in driving avian demographics.
However, several important limitations must be considered when interpreting these results. First, detection biases differ substantially across landscapes. Urban carcasses occur in high-visibility areas and are readily detected, while rural mortality events are often missed because they occur in remote or forested habitats where scavenging and decomposition quickly eliminate evidence. Additionally, our mortality index assumes that the ratio of dead bird observations to total bird observations provides a meaningful proxy for actual mortality rates, but this assumption may not hold if detection probabilities vary systematically across seasons, locations, or species.
3.4 CS4: Resolving Plant Phenology
Figure 10: (Top) Query & Verification. Filtering top INQUIRE-Search outputs for "flowering" phenological stage. (Bottom) Phenological stages occur at significant times throughout the year. (left) density of observations from INQUIRE-Search corresponding to different stages. (right) distribution of observations from INQUIRE-Search in different phenological stages visualized with results from ANOVA/Tukey test.
Results and Analysis
Our analysis shows significant differences in timing for each phenological stage resolved using INQUIRE-Search for common milkweed in southern Quebec. The phenological stage with the best return rate (number of observations matching the queried stage) was flowering (84.5%) followed by seeding (80.5%), senescence (26%), and finally emergence (22.5%). An ANOVA revealed significant differences in stages for the day of year of returned observations, and pairwise differences between stages was tested using a Tukey test. All stages were significantly (p<0.05) different from one another with the exception of emergence and flowering, which showed considerable overlap in timing.
Findings
This case study shows that INQUIRE-Search can be used systematically to resolve detailed phenological information at regional scales for specific species. Some phenological stages, namely, flowering and seeding, were more successfully resolved than emergence and senescence. While this could be a limitation of the models ability to identify relevant observations to those stages (the difference between an emerging plant and a young plant is minor) it is also likely due to bias in the community science observations queried, as observations of flowers and seeds are likely more common as they can aid taxonomic identification [iNaturalist202xHighQualityObservations]. Resolving patterns of senescence will also require specific considerations to account for the fact that senescent plants can also be observed long after senescence, evidenced by the fact we identified observations of senescent milkweed throughout the year. Nonetheless, this sort of fine-scale phenological information would have traditionally required daily or weekly surveys to resolve at the level of a single site, let alone a region like Southern Quebec. Our results demonstrate that INQUIRE-Search has the potential to drastically increase our ability to extract phenological information from the wealth of rapidly growing community science observations.
3.5 C5: Whale Re-Identification
Results and Analysis
Figure 11: (Top) Query & Verification. Filtering top INQUIRE-Search outputs for "white underside of humpback whale fluke" where the fluke was clearly visible and of sufficient quality for pattern matching. (Bottom) Matching individuals to the HappyWhale database and analysis. A deep learning Re-ID model retrieves candidate matches in the HappyWhale database, which we manually verify. We then map repeated sightings of confirmed individuals, demonstrating how INQUIRE-Search can support analyses of movement patterns and habitat use.
Of the 153 candidate images retrieved and verified through INQUIRE-Search, 57 were successfully matched to individuals in the public HappyWhale dataset. Despite inspecting only the top 200 ranked search results, these matches identified 34 unique humpback whales, including multiple sightings of a few individuals, as shown in Fig 11. These additional sightings extend spatiotemporal coverage of known individuals, demonstrating that INQUIRE-Search can uncover valuable, previously untapped data embedded within large community science repositories.
Fig 11 also visualizes matched iNaturalist and HappyWhale observations, highlighting how this linking expands our ability to track individuals across datasets. Since iNaturalist observations include geolocation metadata, cross-referencing matches provides new spatial information for individuals that were previously represented only in curated catalogs. This capability enables researchers to map additional movement paths, identify unusual sightings, and potentially detect shifts in habitat use when opportunistic images align with structured monitoring efforts.
Findings
While small in scale, this case study demonstrates that images can be surfaced by INQUIRE-Search, then linked to known individuals using re-ID, augmenting existing datasets and providing valuable new information for population monitoring. INQUIRE shows strong potential for retrieving high-quality, correctly posed images that capture identifiable features, which enables the integration of valuable data from previously disconnected sources. However, its broader utility is constrained by the identification step, which still depends on a pre-existing labeled dataset and a trained model for matching. Because INQUIRE cannot assign individual identities on its own, its use is more limited when building datasets from scratch or incorporating previously unseen individuals. In addition, retrieval performance reflects the spatial and behavioral biases of community-sourced photographs, meaning the tool will preferentially surface observations from well-sampled areas and charismatic encounters rather than providing an unbiased representation of whale distribution or behavior.
4 Discussion
Across our five case studies, INQUIRE-Search enabled rapid, expert-guided discovery by turning community-sourced images into analyzable ecological data. The system consistently recovered information that would have required substantial fieldwork, helping to resolve major temporal, spatial, and observational gaps.
4.1 Rethinking the Scientific Process in the Age of Image Search
With the interactive, iterative approach to data discovery enabled by INQUIRE-Search, scientists remain the drivers of inquiry, interpretation, and verification. Incorporating such a tool with an ecological research workflow requires rethinking experimental design. The following discussion highlights parameters to "tune" in this process, from identifying suitable questions and crafting precise prompts, to allocating verification effort and accounting for uncertainty.
4.1.1 Which questions are well-posed for INQUIRE-Search?
INQUIRE-Search is fundamentally limited by the images that exist in the underlying repository. Questions targeting extremely rare events, under-observed species, or obscure behaviors may yield insufficient data for robust analysis. For example, in the avian diet case study, species with few iNaturalist records produced too few relevant images of feeding behaviors to support quantitative comparisons across time. Conversely, when data are overly-abundant, as in the wildlife mortality study, search still improves efficiency but does not eliminate the need for substantial verification. Useful queries must also be unambiguously visually verifiable: prompts tied to explicit, observable features (“milkweed with seed pods”) are more reliable than abstract states (“tree under stress”). Well-posed questions therefore target phenomena that are (1) sufficiently represented in the data pool, (2) not so broad or ubiquitous that they produce large volumes results, and (3) visually verifiable.
4.1.2 How to refine a question: iterating on prompts
Search-based discovery depends on clear, specific prompts. Precise language reduces ambiguity in the embedding space and improves the relevance of returned images. Specificity also increases the verifiability of the outputs: "robin pulling a worm from the soil" will yield cleaner results than "bird foraging. Iteration is often necessary, where prompts can be refined by adding behavioral, contextual, or anatomical cues, while overly narrow prompts can be relaxed when returns are sparse.
4.1.3 Managing Effort, Ranking, and Verification
This workflow introduces a new design choice: in addition to formulating effective queries, researchers must also decide how far down the ranked search results to inspect. In this way, "effort" becomes tunable. Reviewing only the top-ranked subset often recovers the most relevant observations with minimal redundancy, as shown in the dietary analysis case study, where experts verified only first 200 images recovered to obtain a sufficient quantity of data.
4.1.4 Identifying Uncertainty and Bias with INQUIRE-Search
INQUIRE-Search inherits the biases of community-science platforms. Observer preferences skew data toward charismatic, accessible, or unusual events, while cryptic interactions and remote habitats remain under-represented [dimson2023observer]. Although methods for observer-effort corrections to quantify and mitigate these biases exist [robinson2018correcting, guilbaultexplorers, mondain2024adaptive, padilla2024assessing], their effectiveness at a global scale remains limited because they require structured sampling, repeated surveys, or explicit effort measurements that community science data often lack.
Model-driven biases add a second layer of uncertainty. VLMs also struggle with prompts requiring complex relational or compositional understanding [alhamoud2025visionlanguagemodelsunderstandnegation, thrush2022winogroundprobingvisionlanguage]. Moreover, recent work shows that VLMs are both miscalibrated (where the expressed confidence does not match true accuracy) [guo2017calibration] and biased (the model defaults to memorized training-set associations instead of analyzing visual evidence) [vo2025vision]. Adapting existing statistical tools for bias estimation and correction [narduzzi2014inverse, geifman2017selective, karimi2023quantifying] could provide a foundation for quantifying uncertainty in discovery-driven search.
4.2 Expanding Beyond iNaturalist
Although our case studies drew on iNaturalist for its scale and ecological breadth, the open-source design of the INQUIRE-Search codebase makes it adaptable to any large image collection. Early qualitative experiments suggest that the system has potential for other ecological image repositories, such as camera trap datasets. Because the system is open-source and resource-efficient, it has the potential to be deployed widely, enabling scientists, NGOs, and local communities to harness image search for ecological monitoring and hypothesis generation on their own databases.
5 Conclusion
Growing ecological image databases remain an underutilized scientific resource. We demonstrate that INQUIRE-Search enables iterative, natural language–driven discovery and verification of diverse ecological signals, supporting scalable analysis of phenomena such as bird diet and plant phenology. This approach suggests a new paradigm in ecological research, in which hypothesis generation, testing, and data curation are integrated within an interactive search system over existing data. By transforming community-sourced imagery from incidental observations into systematic ecological evidence, INQUIRE-Search broadens access to data-driven discovery while raising important questions about rigor, reproducibility, and inference. When combined with careful experimental design, it offers a scalable pathway to accelerate ecological insight in a changing world
Supplementary Information
5.1 System Architecture and Design Motivation
INQUIRE-Search is a tool designed to enable scientists to quickly and easily discover data from within a large ecological database such as iNaturalist. This challenge of surfacing data of interest from a large pool is commonly referred to as "image retrieval" in the AI community, and INQUIRE-Search specifically addresses two fundamental image retrieval challenges to ensure the system is high-performing in terms of both quality and speed. First, INQUIRE-Search enables scientists to search for data using open-ended text, which significantly expands the searchable possibilities beyond existing metadata within the database. Second, INQUIRE-Search is engineered to power this search over very large databases efficiently, without requiring significant computational resources. This efficiency is a key feature of the design, letting scientists iterate on their searches with the system in real-time to discover their data of interest. The system architecture of INQUIRE-Search reflects both of these priorities, combining state-of-the-art vision-language models with efficient indexing and memory management techniques to deliver responsive searches across hundreds of millions of images on modest hardware (4 vCPUs with 32GB RAM, roughly the capacity of a single laptop).
5.2 Search Index
Searching for relevant data using text is accomplished by first embedding the text query using the same SigLIP model used to embed the iNaturalist images. We then calculate the vector similarity between that new text embedding and each image embedding and use that similarity score to rank the full set of images to surface possible images of interest. With a smaller database, this entire process can be accomplished directly within the memory of a modest computer, but when scaling to the hundreds of millions of images in iNaturalist the set of image embeddings becomes large ($\sim$200GB), requiring a different approach to quickly find the most similar images to a new text query.
Vector databases, which use a specialized indexing system that allows fast search over millions of high-dimensional vectors, [decastro2018detecting, singla2021experimental, fang2017topology] are specifically designed to facilitate efficient similarity-based search. We utilize FAISS [douze2025faiss] to create an approximate nearest neighbor index that is small and fast to query. Our FAISS index is generated and tuned using the Autofaiss library, a widely-used library for efficient nearest-neighbor search [webster2023duplication]. The final index is memory-mapped (stored on disk but accessed as if it were in memory, enabling fast search with limited RAM), uses 36GB of storage, and enables sub-500ms search times while maintaining high retrieval accuracy.
In addition to the visual embedding index, our search tool supports several metadata-based mechanisms to filter by:
  1. 1.
    Taxonomy: We map each taxon (species, genus, family, etc.) to a list of image IDs known to contain that entity, using taxon labels provided for each image from iNaturalist. This allows for efficient subset selection when users apply taxonomic filters to their queries.
  2. 2.
    Temporal Range: Users can filter search queries by month.
  3. 3.
    Geographic Extent: Using approximate image location data, optionally provided by iNaturalist contributors, users can filter images to a specific geographic area using latitude-longitude bounds.
5.3 Image and Text Embeddings
The core search capability behind INQUIRE-Search is powered by such a vision-language model. Not all vision-language models are equally capable of ranking ecological data, particularly for scientific queries. We leveraged our previous work developing the INQUIRE-Benchmark, which specifically evaluates models for ecological image retrieval tasks, to select SigLIP-So400m-384-14 [zhai2023sigmoid] as our embedding model, due to both its strong performance on the benchmark and its reasonable tradeoff between speed and accuracy; other models like ViT-H-14 [radford2021learning] are slightly more accurate at retrieving relevant data but are significantly larger, and thus slower. We processed 300M images sourced from iNaturalist through the visual encoder component of the selected SigLIP model to build the backbone of the INQUIRE-Search tool.
Figure 12: INQUIRE-Search user interface. INQUIRE-Search provides an user-friendly and intuitive interface that allows ecologists to query hundreds of millions of natural-world images. This example illustrates a query for "American robin eating," with species and temporal filters applied, enabling rapid retrieval, inspection, and verification of relevant observations.
5.4 Search Interface
INQUIRE-Search provides a web-based interface (Fig 12) that surfaces the search index and filters to scientists. The interface includes a query field for natural language queries (e.g., "California condor with a green ‘26’ tag on its wing") and controls for taxonomic, temporal, or geospatial filters. The interface displays a grid of thumbnails of the images returned from the user-provided query, ordered by relevance to the query. The images are clickable, allowing the user to mark images for further analysis. An expanded view of a selected image shows a full-resolution image with complete metadata.
5.5 Data Export
When the user exports the data, a CSV file is prepared for download. This file contains a row for each image which includes a field for whether an image was marked by the user, detailed metadata, and a link to the associated iNaturalist observation.
5.6 Data Analysis
Once verified, these retrieved observations form the foundation for a wide range of statistical analyses, demonstrated in the following section. This workflow allows experts to effectively "design" their search strategy and specify what ecological signals to target, while identifying or mitigating relevant sources of uncertainty and bias.
5.7 Retrieval and Filtering Results
Species Prompt Template Season Inspected Filtered Rate
(Inv/Vert/Seed/Fruit/Nect/Carr/Plant)
American Robin "<species> with <diet type> in its mouth" Summer 500 per type 338 67.6%
Winter 500 per type 256 51.2%
Red-bellied Woodpecker Summer 500 per type 95 19.0%
Winter 500 per type 161 32.2%
American Tree Sparrow Summer 500 per type 1 0.2%
Winter 500 per type 83 16.6%
Ancient Murrelet Summer 500 per type 0 0.0%
Winter 500 per type 2 0.4%
Gray-cheeked Thrush Summer 500 per type 0 0.0%
Winter 500 per type 0 0.0%
Key: Inv=Invertebrate, Vert=Vertebrate, Nect=Nectar, Carr=Carrion. Counts are placeholders (-/…) where detailed breakdowns were not provided in source text. Total inspected per diet/season/species is 500.
Table 2: Retrieval results for seasonal variation in bird diets. The prompt template is applied to all species. Diet types are abbreviated in the header row.
Target Prompt Filters Insp. Filt. Rate
Young coniferous trees "young coniferous trees in burned forest" Geo: High Park Fire (40.57-40.75 N, 105.18-105.54 W); Date: Post-2012 100 78 78.0%
Young deciduous trees "young deciduous trees in burned forest" 112 45 40.2%
Filters sourced from Table 2.
Table 3: Search parameters and results for post-fire forest regrowth.
Location Prompt Filters Insp. Filt. Rate
Boston (Urban) "dead bird" Geo: Bounding box around Boston, MA; No Taxon filter 1000 543 54.3%
Pioneer Valley (Rural) "dead bird" Geo: Bounding box around Pioneer Valley/Amherst, MA; No Taxon filter 360 79 21.9%
Table 4: Search parameters and results for wildlife mortality.
Stage Prompt Filters Insp. Filt. Rate
Emergence "Milkweed germinating or emerging" Taxon: Asclepias syriaca; Geo: S. Quebec (45.03 to 46.54 N, -74.68 to -71.66 W); No temporal filter 200 45 22.5%
Flowering "Milkweed flowering" 200 169 84.5%
Seeding "Milkweed producing seeds…" 200 161 80.5%
Senescence "Milkweed dying or withering…" 200 52 26.0%
Table 5: Search parameters and results for plant phenology stages.
Target Prompt Filters Insp. Filt. Rate
Humpback whale flukes "white underside of humpback whale fluke" Taxon: Humpback Whale (Megaptera novaeangliae); No Geo/Temp filters 200 153 76.5%
Table 6: Search parameters and results for humpback whale re-identification.